OpenStudy (anonymous):

The dimension of triangle A are twice the dimension of triangle B..The area of trigle A is 98cm2....................WHta is the area of tringle B

jhonyy9 (jhonyy9):

area = base * height /2 A area = base *height /2 = 98 cm^2 B area = 2base *2height /2 = 2*base*height = 4*98 =392 cm^2

Directrix (directrix):

By the SSS Similarity Theorem, Triangle A is similar to Triangle B. The scale factor from A to B is 2 to 2 or 2/1. If two triangles are similar, the square of their scale factor is equal to the ratio of their areas. (2/1)^2 = 98/B where B represents the area of triangle B 4/1 = 98/B 4 B = 98 B = 98/4 = 49/2 square cm